zum Inhalt springen

News-Archiv

22.2.2019 - NMR-Abteilung: Server operating, change in N2 refill schedule (carnival)

After renovation, our LIMS and database server operates again. During carnival, there are changes for the N2 refill…

14.2.2019 - NMR-Abteilung: AV300oc schedule changes (OC-WP) until March 15th

Currently, the AV300oc spectrometer has a changed schedule for walk-by operation due to Praktikum users

13.2.2019 - NMR-Abteilung: AV300oc access restricted in the afternoon

Today, students of the organic chemistry practical course will be trained at the AV300oc during afternoon

11.2.2019 - NMR-Abteilung: AV499 service interrupted tomorrow

There will be a revision in our pressurized air system tomorrow morning.

4.2.2019 - NMR-Abteilung: Labs closed for cleaning

On Friday, 22th of February the floor of NMR facility labs 115 and 116 will be cleaned and may not be entered during…

15.1.2019 - NMR-Abteilung: Network renovation on February 19th

On 19 February the NMR facility network will move into a new subnet. This will disturb and partially interrupt the…

7.1.2019 - NMR-Abteilung: Nutzerinformationen fuer 2019 / first circular user mail 2019

Wir wuenschen Ihnen ein Gutes Neues Jahr und hoffen, dass sich die seit dem vergangenen Sommer begonnenen…

3.1.2019 - NMR-Abteilung: AV300oc back in service

After cleaning of the heavily polluted RF and shim unit, the spectrometer is back in service. Please check the…

02.01.2019 - NMR-Abteilung: Downtime AV300oc tomorrow

Tomorrow (January, 3rd) the av300oc will be out of service due to necessity of a probe and shim stack cleaning.

20.12.2018 - Schließung der Abteilung während der Feiertage

Während der Weihnachtszeit bleibt die NMR-Abteilung vom 21.12.18 13:00 Uhr bis einschließlich 1.1.2019 geschlossen.

7.12.2018 - NMR-Abteilung: av300oc

The av300oc started with some difficulties but is now working fine. The overnight measurements are now running until…

4.12.2018 - NMR-Abteilung: AV300ac out of service

Leider ist heute ein neues Problem an der Automation des AV300ac aufgetreten. Das Spektrometer muss daher bis auf…

30.11.2018: AV300oc & AV300ac / AV499

The AV300oc unfortunately will be out service for a couple of days. There is a hardware damage which needs to be…

30.11.18 - NMR-Abteilung: Automation disrupted

Leider ist die Automation am av300oc gestern Abend abgebrochen. Eine Probe ist nicht hochgekommen, was *wieder einmal*…

28.11.2018 - Temporary downtime tomorrow morning, AV500 renovation / Zeitweise Abschaltung

Due to work affecting the electric circuits on our floor, tomorrow morning, temporary power failure will lead to…

zum Seitenanfang